# Language Models Do Not Follow Occam’s Razor: A Benchmark for Inductive and Abductive Reasoning

### v1

@@ Appendix C Algorithms
 generates the observations where selected members also belong to a supertype
end if
@@
return all world model axioms, observations, and the ground truth hypothesis
Algorithm 1 Generate a single hypothesis
// First, build the ontology tree
// $Node()$ will initialize a node object
// Add a root node
$tree\_nodes\leftarrow[[Node()]]$
for $i$ from 1 to h1h-1 do
$current\_nodes\leftarrow[]$
for $parrent\_node\in tree\_nodes[-1]$ do
  $u\leftarrow\texttt{random}(0,1)$
  if $u>0.5$ then
   for $j$ from 1 to 2 do
    current_nodes.append(Node())current\_nodes.append(Node())
   end for
  else
   for $j$ from 1 to 3 do
    current_nodes.append(Node())current\_nodes.append(Node())
   end for
  end if
end for
tree_nodes.append(current_nodes)tree\_nodes.append(current\_nodes)
end for
for $current\_nodes\in tree\_nodes$ do
for $node\in current\_nodes$ do
  // Allocate three members and properties
  for $j$ from 1 to 3 do
   $node.members.append(RandomMember())$
   $node.properties.append(RandomProperty())$
  end for
end for
end for
if single hypothesis then
 Randomly select a task type
 generates a single hypothesis
else
 generate 1 to 3 hypothesis per tree layer
1end if
Algorithm 2 Generate a reason example
Appendix D More results
@@
Appendix E LLM example responses
Self-verification
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
Hypotheses: Dalpists are rainy.
LLM Output: <think> … Suppose Dalpists are rainy. Since Amy and Jerry and Pamela are dalpists, they are all rainy. This hypothesis can explain all observations. </think>
Wrong ontology direction
@@ Appendix E LLM example responses
World Model: Porpors are dalpists. Nicole is a gergit. Sarpers are dalpists. Every gergit is a dalpist. Debra is a dalpist. Each dalpist is rainy. Every porpor is not cold. Raymond is a gergit. Each porpor is translucent. Amy is a porpor. Jerry is a dalpist. Pamela is a dalpist. Sharon is a gergit. Edward is a dalpist. Michael is a sarper.
Observations: Sharon is moderate. Barbara is not cold. Debra is liquid. Barbara is rainy. Barbara is liquid. Barbara is translucent. Amy is a rompus. Raymond is moderate. Jerry is a rompus. Edward is liquid. Pamela is a rompus. Nicole is moderate.
Ground truth hypotheses: Dalpists are rompuses. Every dalpist is liquid. Each gergit is moderate. Barbara is a porpor.
LLM Output: … Rompuses are dalpists. …
Ignore the ontology and produce unnecessary hypotheses
@@
World Model: David is a sorple. Each sorple is a dropant. Each hilper is a thorpin. Rimpees are scrompists. Brian is a sorple. Dropants are gompers. Angela is a gomper. Each thorpin is a gomper. Every scrompist is a gomper.
Observations: David is rainy. Brian is rainy. Angela is rainy.
Ground truth hypotheses: Every gomper is rainy.
LLM Output: … Every gomper is rainy. Every dropant is rainy. …
Fall back to rival hypothesee
@@
World Model: Rorpants are dolpers. Every worple is a wumpus. Each drompee is moderate. Ashley is a rorpant. Fimples are dolpers. Ronald is a wumpus. Wumpuses are gwompants. Yompins are dolpers. Brian is a dolper. Michelle is a worple. Each dolper is a gwompant. Shirley is a rorpant. Sandra is a bempin. Linda is a fimple. Drompees are pale. Jack is a yompin. Each drompee is a bempin. Bempins are cold. Bempins are gwompants. Emma is a gwompant. Michael is a gwompant.
Observations: Shirley is a borpin. Christopher is pale. Emma is a borpin. Brian is salty. Michael is salty. Christopher is a borpin. Christopher is cold. Ashley is salty. Christopher is moderate.
Ground truth hypotheses: Each gwompant is a borpin. Each gwompant is salty. Christopher is a drompee.
LLM Output: … Shirley is a borpin. …
Hallucinated entities
@@
World Model: Jonathan is a rifpist. Yimples are dull. Samantha is a rifpist. Karen is a gwompant. Scott is a jelgit. Olivia is a frompor. Maria is a lirpin. Twimpees are lirpins. Each rifpist is aggressive. Frompors are jelgits. Jessica is a jelgit. Amanda is a twimpee. Rifpists are jelgits. Each lirpin is a jelgit. Yimples are small. Each gwompant is a frompor. Each yimple is a rifpist.
Observations: Mary is not overcast. Scott is not overcast. Jonathan is not overcast. Samantha is a kurpor. Mary is aggressive. Jessica is a kurpor. Olivia is a kurpor. Mary is small. Mary is dull.
Ground truth hypotheses: Every jelgit is a kurpor. Jelgits are not overcast. Mary is a yimple.
LLM Output: … Mary is a dolphant. …
Confuse members with concepts
World Model: Delpees are impuses. Yumpuses are delpees. Rebecca is an impus. Kathleen is a rifpist. Each impus is rainy. Thomas is a delpee. Lompees are arpers. Each bempin is an impus. Stephen is a bempin. Every impus is shy. Laura is a yumpus. Every arper is an impus. Stephanie is an arper. Lompees are moderate. Rifpists are bempins.
Observations: Frank is moderate. Frank is rainy. Frank is shy.
Ground truth hypotheses: Frank is a lompee.
LLM Output: … Lompee is Frank. …

### v2

@@ Appendix F Algorithms
18       generate the observations where selected members also belong to a supertype
19
@@
20   return all world model axioms, observations, and the ground-truth hypothesis
Algorithm 1 Generate a single hypothesis.
1function generate_example()
2    /* first, build the ontology tree */
3    /* node() initializes an empty node object */
4    /* add a root node */
5    $tree\_nodes\leftarrow$ [[node()]]
6    for $i=1,\ldots,h-1$ do
7       $current\_nodes\leftarrow[]$
8       for $parent\_node\in tree\_nodes[-1]$ do
9          $u\leftarrow$ random(0,1)
10          if $u>\frac{1}{2}$
11             repeat 2 times
12                current_nodescurrent\_nodes.append(node())
13
14
15         else
16             repeat 3 times
17                current_nodescurrent\_nodes.append(node())
18
19
20
21      tree_nodestree\_nodes.append(current_nodescurrent\_nodes)
22
23   for $current\_nodes\in tree\_nodes$ do
24       for $node\in current\_nodes$ do
25          /* allocate three members and properties */
26          repeat 3 times
27             nodenode.members .append( random_member())
28             nodenode.properties .append( random_property())
29
30
31
32   if this is a single-hypothesis example
33       randomly select a task type
34       generate_hypothesis(tree_nodestree\_nodes)
35
36   else generate 1 to 3 hypotheses per tree layer
37
Algorithm 2 Generate an abductive/inductive reasoning example.
Appendix G Results with absolute accuracy
@@
Appendix H Error analysis
H.1 Common error types and example responses
In this section, we detail the most common types of errors that LLMs produce in our experiments. For each of the error types listed below, we provide an example consisting of a world model, list of observations, ground-truth hypotheses, and the mistaken LLM output.
@@ H.1 Common error types and example responses
Error Type 1: Wrong ontology direction World Model: Porpors are dalpists. Nicole is a gergit. Sarpers are dalpists. Every gergit is a dalpist. Debra is a dalpist. Each dalpist is rainy. Every porpor is not cold. Raymond is a gergit. Each porpor is translucent. Amy is a porpor. Jerry is a dalpist. Pamela is a dalpist. Sharon is a gergit. Edward is a dalpist. Michael is a sarper. Observations: Sharon is moderate. Barbara is not cold. Debra is liquid. Barbara is rainy. Barbara is liquid. Barbara is translucent. Amy is a rompus. Raymond is moderate. Jerry is a rompus. Edward is liquid. Pamela is a rompus. Nicole is moderate. Ground-truth hypotheses: Dalpists are rompuses. Every dalpist is liquid. Each gergit is moderate. Barbara is a porpor. Mistaken LLM Output: … Rompuses are dalpists. …
@@
Error Type 2: Ignore the ontology and produce unnecessary hypotheses World Model: David is a sorple. Each sorple is a dropant. Each hilper is a thorpin. Rimpees are scrompists. Brian is a sorple. Dropants are gompers. Angela is a gomper. Each thorpin is a gomper. Every scrompist is a gomper. Observations: David is rainy. Brian is rainy. Angela is rainy. Ground-truth hypotheses: Every gomper is rainy. Mistaken LLM Output: … Every gomper is rainy. Every dropant is rainy. …
@@
Error Type 3: Fall back to rival hypotheses World Model: Rorpants are dolpers. Every worple is a wumpus. Each drompee is moderate. Ashley is a rorpant. Fimples are dolpers. Ronald is a wumpus. Wumpuses are gwompants. Yompins are dolpers. Brian is a dolper. Michelle is a worple. Each dolper is a gwompant. Shirley is a rorpant. Sandra is a bempin. Linda is a fimple. Drompees are pale. Jack is a yompin. Each drompee is a bempin. Bempins are cold. Bempins are gwompants. Emma is a gwompant. Michael is a gwompant. Observations: Shirley is a borpin. Christopher is pale. Emma is a borpin. Brian is salty. Michael is salty. Christopher is a borpin. Christopher is cold. Ashley is salty. Christopher is moderate. Ground-truth hypotheses: Each gwompant is a borpin. Each gwompant is salty. Christopher is a drompee. Mistaken LLM Output: … Shirley is a borpin. …
@@
Error Type 4: Hallucinated entities World Model: Jonathan is a rifpist. Yimples are dull. Samantha is a rifpist. Karen is a gwompant. Scott is a jelgit. Olivia is a frompor. Maria is a lirpin. Twimpees are lirpins. Each rifpist is aggressive. Frompors are jelgits. Jessica is a jelgit. Amanda is a twimpee. Rifpists are jelgits. Each lirpin is a jelgit. Yimples are small. Each gwompant is a frompor. Each yimple is a rifpist. Observations: Mary is not overcast. Scott is not overcast. Jonathan is not overcast. Samantha is a kurpor. Mary is aggressive. Jessica is a kurpor. Olivia is a kurpor. Mary is small. Mary is dull. Ground-truth hypotheses: Every jelgit is a kurpor. Jelgits are not overcast. Mary is a yimple. Mistaken LLM Output: … Mary is a dolphant. …
H.2 Details of error analysis
We first randomly sampled 200 LLM example responses to reasoning questions of varying difficulty. Then, we manually inspected them to identify recurring failure patterns as summarized in the paper. To further quantitatively study the error patterns of all responses, we use an LLM-as-a-judge (llmasjudge) and prompt GPT-4o with in-context examples. We use the following prompt in Figure 10 for the LLM-as-a-judge. We categorize the error pattern distribution by ontology tree height and present the results in Figure 9. As shown in the figure, Error Type 2 (Ignore the ontology and produce unnecessary hypotheses) is the most significant error across all ontology tree heights, indicating that LLMs can generate weakly correct hypotheses but fail to consider the given ontology tree and optimize them in the sense of Occam’s razor.
Figure 9: Distribution of error types across ontology tree heights.
Prompt for Error Analysis Prompt Template You are an expert evaluator of reasoning quality. Your task is to analyze a model’s reasoning process for a question and determine the type of error it contains. ## Task Given: * A question
* The model’s reasoning process
* The correct answer
Identify which error type best describes the mistake in the reasoning. Choose exactly one error type from the following categories. --- ## Error Types ** Error Type 1: Wrong ontology direction** The hypothesis contains the wrong ontology direction. ** Error Type 2: Fall back to trivial hypotheses** The hypothesis is the observation itself. ** Error Type 3: Ignore the ontology and produce unnecessary hypotheses** The hypothesis is technically correct but unnecessary given the ontology. ** Error Type 4: Hallucinated entities** The reasoning relies on non-existent entities. --- ## Examples ### Example 1 (Wrong ontology direction) <omitted for space, see §H.1 for examples> --- ### Example 2 (Fall back to trivial hypotheses) <omitted for space, see §H.1 for examples> --- ### Example 3 (Ignore the ontology and produce unnecessary hypotheses) <omitted for space, see §H.1 for examples> --- ### Example 4 (Hallucinated entities) <omitted for space, see §H.1 for examples> --- ## Now evaluate the following case Question: {{QUESTION}} Model Reasoning: {{MODEL_REASONING}} Correct Answer: {{CORRECT_ANSWER}} --- ## Output Format Error Type: <one of: Wrong ontology direction | Fall back to trivial hypotheses | Ignore the ontology and produce unnecessary hypotheses | Hallucinated entities>
Figure 10: Prompt used to analyze common error patterns.